OpenStudy (anonymous):

A rubber ball dropped on a hard surface takes a sequence of bounces, each one 3/5 as high as the preceding one. If this ball is dropped from a height of 10 feet, what is the total vertical distance it has traveled after it hits the surface the 5th time?

OpenStudy (anonymous):

@Callisto lol sorry i got ur name wrong last time

OpenStudy (anonymous):

@adjaharrison

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

Well, i hate word problems :P lemme try this out :)

OpenStudy (callisto):

It's sum of 2 GS ...

OpenStudy (callisto):

For the first GS: a =10; r = 3/5 For the second GS: a = 10(3/5) = 6 ; r =3/5 Use the formula Sum = a/(1-r) add them , like this Total distance = 10 / [1 - (3/5)] + 6 / [1-(3/5)] =?